# Young Diagram Decompositions for Almost Symmetric Numerical Semigroups

### v1

@@ 1. Introduction
Throughout, we denote by $\mathbb{Z}$ and $\mathbb{N}$ the sets of integers and positive integers, respectively, and define $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. A numerical set $R$ is a subset of $\mathbb{N}_{0}$ that contains 00 and has a finite complement in $\mathbb{N}_{0}$. In particular, $\mathbb{N}_{0}$ itself is a numerical set with an empty complement. If a numerical set $R$ is closed under addition, i.e. if $x,y\in R$ implies $x+y\in R$, then it is referred to as a numerical semigroup.
@@ 2. Numerical Sets and Semigroups
If a numerical set $R$ is not equal to $\mathbb{N}_{0}$, it is referred to as proper. For a proper numerical set $R$, the complement $\operatorname{G}(R)=\mathbb{N}_{0}\setminus R$ is called the set of gaps of $R$. The genus of $R$, denoted by $\operatorname{g}(R)$, is the number of gaps of $R$, i.e. the cardinality of $\operatorname{G}(R)$. The largest element of $\operatorname{G}(R)$ is known as the Frobenius number of $R$, denoted by $\operatorname{F}(R)$. The quantity $\operatorname{C}(R)=\operatorname{F}(R)+1$ is referred to as the conductor of $R$, which satisfies that $z\in\mathbb{N}_{0}$ and $z>\operatorname{C}(R)$ implies $z\in R$. Also, we define the Frobenius of $\mathbb{N}_{0}$ to be $-1$, and its conductor of to be 00.
@@
A numerical set $R$ is said to be symmetric if for all $a\in[0,\operatorname{F}(R)]\cap\mathbb{Z}$, exactly one of the elements $a$ or $\operatorname{F}(R)-a$ belongs to $R$.
@@
A numerical semigroup $R$ is symmetric if it is symmetric as a numerical set, or equivalently if it has the property that $12\operatorname{g}(R)=\operatorname{F}(R)+$ 1. And $R$ is called pseudo-symmetric if and only if $\operatorname{F}(R)$ is even and for every $z\in\mathbb{Z}\setminus R$, either $z\in\{\frac{\operatorname{F}(R)}{2},\operatorname{F}(R)\}$ or $\operatorname{F}(R)-z\in H$. Equivalently $R$ is pseudo-symmetric if and only if $22\operatorname{g}(R)=\operatorname{F}(R)+$ 2.
@@ 3. Numerical Sets and Young Diagrams
Let $Y$ be a Young diagram with $n$ columns and $g$ rows. The number of boxes in a given column (respectively, row) is referred to as the length of that column (respectively, row). The hook of a box in $Y$ consists of the box itself, all boxes to its right in the same row, and all boxes below it in the same column. The total number of boxes in the hook is called the hook length of that box.
@@
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{\longleftrightarrow}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{\longleftrightarrow}$
Let $R=\{0,r_{1},\dots,r_{n-1},r_{n},\rightarrow\}$ be a numerical set with gap set $\operatorname{G}(R)=\{a_{1},\dots,a_{g}\}$ and associated Young diagram $Y_{R}$. By construction, $Y_{R}$ has $g$ rows and $n$ columns. From left to right, for each $i$th small element of $R$ corresponds to the $(i+1)$th column of $Y_{R}$, so that $r_{0}$ corresponds to the first column, $r_{1}$ to the second column, and so on. Similarly, from bottom to top, for each gap $a_{j}$ of $R$ corresponds to the $j$th row of $Y_{R}$.
The hook length of the box in the $i$th row (from bottom to top) and the first column of $Y_{R}$ is the $i$th gap $a_{i}$ of $R$. Moreover, for each $i\in[0,n-1]\cap\mathbb{Z}$, the $i$th column of $Y_{R}$ can be identified with the set of hook lengths of the boxes in it. We denote this set by $\operatorname{G}_{i}(R)$.
@@
[15, Section 2] Let $R=\{0,r_{1},\dots,r_{n-1},r_{n},\rightarrow\}$ be a numerical set with the associated Young diagram $Y_{R}$. Then:
@@
    For each $i\in[0,n-1]\cap\mathbb{Z}$, we have
    $\operatorname{G}_{i}(R)=\{a-r_{i}\mid a\in G(R)\text{ and }a>r_{i}\}=\operatorname{G}(R-r_{i})$
    where $R-r_{i}=\{0,r_{i+1}-r_{i},\dots,r_{n}-r_{i},\rightarrow\}$.
@@
    $R$ is a numerical semigroup if and only if $\operatorname{G}_{i}(R)\subseteq\operatorname{G}_{0}(R)$ for all $i\in[0,n-1]\cap\mathbb{Z}$, or equivalently $\bigcup_{i=0}^{n-1}\operatorname{G}_{i}(R)=\operatorname{G}(R)$.
By Proposition 3.2, it follows that $\operatorname{G}_{0}(R)=\operatorname{G}(R)$, and for each $i\in[0,n-1]\cap\mathbb{Z}$, the hook length of the box in the $i$th column (from left to right) and the top row of $Y_{R}$ is $\operatorname{F}(R)-r_{i}$.
@@
For a positive integer $N$, a partition $\nu$ of $N$ is a finite non-increasing sequence of positive integers $\nu_{1}\geq\nu_{2}\geq\dots\geq\nu_{n}$ such that $\nu_{1}+\nu_{2}+\dots+\nu_{n}=N$, denoted by $\nu=(\nu_{1},\nu_{2},\dots,\nu_{n})$. Each $\nu_{i}$ is called a part of the partition and the number $n$ of parts is referred to as the length of the partition.
In the context of Young diagrams, listing all the lengths of all columns from left to right yields a partition. Conversely, every partition $\nu=[\nu_{1},\nu_{2},\dots,\nu_{n}]$ corresponds to a Young diagram with $\nu_{1}$ columns and $n$ rows, where the length of $i$th row is $\nu_{i}$. This correspondence defines a bijective mapping between the set of partitions and the set of Young diagrams. Hence, we can identify a Young diagram $Y$ as $Y=\nu$ if its corresponding partition is $\nu$. For instance, the associated Young diagram $Y_{R}$ of the numerical set $R$ given in 2.1 is corresponding to the partition $[6,4,2,1,1,1]$ of $15$, and so we write $Y_{R}=[6,4,2,1,1,1]$.
Next, we present algorithms for visualizing numerical sets using Young diagrams. To facilitate their description, we begin by establishing some notation. Let $L=[l_{1},\dots,l_{n}]$ be a list of length $n$. We denote by $L[i]$ the $i$th element of $L$. For an integer $a$, we define $a\pm L$ as the list $[a\pm l_{1},\dots,a\pm l_{n}]$. Additionally, $\operatorname{Filtered}(L,m\rightarrow m>0)$ denotes the list containing only the positive elements of $L$, preserving their original order.
@@
0:  A list of integers starting from zero in increasing order $[r_{0}=0,r_{1},\dots,r_{n}]$.
@@
0:  A list of non-increasing positive integers $[p_{1},\dots,p_{n}]$.
@@
0:  A list of integers starting from zero in increasing order $[r_{0}=0,r_{1},\dots,r_{n}]$.
@@ 4. Operations on Young Diagrams
[…]
$\underbrace{\hskip 24.00002pt}_{\text{first case}}$ $\underbrace{\hskip 24.00002pt}_{\text{second case}}$ $\underbrace{\hskip 24.00002pt}_{\text{third case}}$
@@
    The bonded sum of $Y$ and $X$, $Y\boxplus_{B}X$, is defined by glueing $X$ above $Y$ in such a way that the bottom row of $X$ is placed to right of the top row of $Y$. The resulting diagram is a Young diagram with $n+m$ columns and $g+k-1$ rows.
@@
    The conjoint sum of $Y$ and $X$, $Y\boxplus_{C}X$, is defined by glueing $X$ above $Y$ such that the most-left column of $X$ is placed directly on top of the right-most column of $Y$. The resulting diagram is a Young diagram with $n+m-1$ columns and $g+k$ rows.
@@
[…]
    $\underbrace{\hskip 31.99997pt}_{\displaystyle Y}$ $\underbrace{\hskip 31.99997pt}_{\displaystyle X}$ $\underbrace{\hskip 64.00002pt}_{\displaystyle Y\boxplus_{B}X}$
@@
    $\underbrace{\hskip 31.99997pt}_{\displaystyle Y}$ $\underbrace{\hskip 31.99997pt}_{\displaystyle X}$ $\underbrace{\hskip 64.00002pt}_{\displaystyle Y\boxplus_{E}X}$
@@
[…]
    $\underbrace{\hskip 31.99997pt}_{\displaystyle Y}$ $\underbrace{\hskip 31.99997pt}_{\displaystyle X}$ $\underbrace{\hskip 55.99998pt}_{\displaystyle Y\boxplus_{C}X}$
@@
Over and above that we get the conductor of the resulting numerical set $R\boxplus_{C}S$ as $\operatorname{C}(R\boxplus_{C}S)=\operatorname{C}(R)+\operatorname{C}(S)-1$, and the Frobenius of $R\boxplus_{C}S$ as $\operatorname{F}(R\boxplus_{C}S)=F(R)+F(S)$. By analogous descriptions conjoint and end-to-end sums of $R$ and $S$ can be formulated as well.
@@
Let $R=\{0,r_{1},\dots,r_{n}=\operatorname{C}(R),\rightarrow\}$ and $S=\{0,s_{1},\dots,s_{m}=\operatorname{C}(S),\rightarrow\}$ be proper numerical sets. Then we define
@@
Let $R=\{0,r_{1},\dots,r_{n},\rightarrow\}$ and $S=\{0,s_{1},\dots,s_{m},\rightarrow\}$ be proper numerical sets with $\operatorname{G}(R)=\{a_{1},\dots,a_{g}\}$ and $\operatorname{G}(S)=\{b_{1},\dots,b_{k}\}$, respectively. Then
@@
    $\operatorname{G}(R\boxplus_{B}S)=\{a_{1},\dots,a_{g-1},a_{g}+b_{1},\dots,a_{g}+b_{k}\}$ and $\operatorname{F}(R\boxplus_{B}R)=a_{g}+b_{k}$,
@@
Let $R=\{0,r_{1},\dots,r_{n},\rightarrow\}$ be a proper numerical set. Then $\mathbb{N}_{0}\boxplus_{B}R=\{-1,0,r_{1}-1,\dots,r_{n}-1,\rightarrow\}$ is not a numerical set, and $\mathbb{N}_{0}\boxplus_{C}R=\{r_{1}-1,\dots,r_{n}-1,\rightarrow\}$ is not also a numerical set unless $r_{1}=1$. However, we naturally have the following.
@@
Let $R$ and $S$ be a numerical sets. Then
@@
Let $R=\{0,r_{1},\dots,r_{n},\rightarrow\}$ with $\operatorname{G}(R)=\{a_{1},\dots,a_{g}\}$ and $S=\{0,s_{1},\dots,s_{m}\}$ with, $\operatorname{G}(S)=\{b_{1},\dots,b_{k}\}$. Then
@@ 5. Decompositions of Almost Symmetric Numerical Semigroups
$R=S\boxplus_{C}\{0,2,\rightarrow\}\boxplus_{B}S^{*}\text{ or }R=S\boxplus_{B}\{0,2,\rightarrow\}\boxplus_{C}S^{*}.$
@@
In case $\frac{C(R)}{2}\notin R$, we select $S=\{0,r_{1},\dots,r_{k-1},r_{k}+1,\rightarrow\}$ with $\operatorname{G}(S)=\{a_{1},\dots,a_{n-k},r_{k}\}$. Then we get $F(R)-\frac{\operatorname{C}(R)}{2}\in R\implies\frac{\operatorname{C}(R)}{2}-1\in R\implies\operatorname{F}(S)=r_{k}=\frac{\operatorname{C}(R)}{2}-1$ and $\operatorname{C}(S)=r_{k}+1=\frac{\operatorname{C}(R)}{2}$. The dual of $S$ is then $S^{*}=\{0,r_{k}-a_{n-k-1},\dots,r_{k}-a_{1},r_{k}+1,\rightarrow\}.$ Therefore,
@@
[…]
$\underbrace{\hskip 15.99998pt}_{\displaystyle Y_{S}}$ $\underbrace{\hskip 8.00003pt}_{\displaystyle Y_{\{0,2,\rightarrow\}}}$ $\underbrace{\hskip 24.00002pt}_{\displaystyle Y_{S^{*}}}$ $\underbrace{\hskip 40.0pt}_{\displaystyle Y_{R}=Y_{S}\boxplus_{C}Y_{\{0,2,\rightarrow\}}\boxplus_{B}Y_{S^{*}}}$
@@
[…]
$\underbrace{\hskip 15.99998pt}_{\displaystyle Y_{S}}$ $\underbrace{\hskip 8.00003pt}_{\displaystyle Y_{\{0,2,\rightarrow\}}}$ $\underbrace{\hskip 31.99997pt}_{\displaystyle Y_{S^{*}}}$ $\underbrace{\hskip 48.00003pt}_{\displaystyle Y_{R}=Y_{S}\boxplus_{B}Y_{\{0,2,\rightarrow\}}\boxplus_{C}Y_{S^{*}}}$
When decomposing a symmetric numerical semigroup $R=\{0,r_{1}\dots,r_{n},\rightarrow\}$ into a numerical semigroup and its dual, some certain considerations must be made. By Theorem 5.1, when $\frac{\operatorname{C}(R)}{2}\notin R$, the natural choice would be $S=\{0,r_{1},\dots,r_{k-1},r_{k}+1,\rightarrow\}$ where $r_{k}=\frac{\operatorname{C}(R)}{2}-1$ for some $k\in[1,n]\cap\mathbb{Z}$. In this case, if $r_{k}$ is not a minimal generator of $R$, then $S$ fails to be a numerical semigroup. Thus, to ensure $S$ remains a numerical semigroup, we need to consider a slightly different $S$ and a different sum.
Hence, when $R=\{0,r_{1}\dots,r_{n},\rightarrow\}$ is a symmetric numerical semigroup with $\frac{\operatorname{C}(R)}{2}\notin R$, to ensure $S$ remains a numerical semigroup in the decomposition, we need to choose that $S=\{0,r_{1},\dots,r_{k},r_{k}+2\rightarrow\}$ with $r_{k}=\frac{\operatorname{C}(R)}{2}-1$ for some $k\in[1,n]\cap\mathbb{Z}$. In this way, we guarantee that $S$ is a numerical semigroup. However, the decomposition takes the form $R=S\boxplus_{C}(S^{*}-1)$ where $S^{*}-1=\{s-1\mid s\in S^{*}\cap\mathbb{N}\}\cup\{0\}$. To prove this statement, we shall use a similar way as in the proof of Theorem 5.1.
@@
[12, cf. Theorem 4.4] Let $R$ be a symmetric numerical semigroup. Then there exists a unique numerical semigroup $S$ such that
@@
Let $R$ be a numerical semigroup with type $\operatorname{t}(R)=t$, and let $\operatorname{PF}(R)=\{a_{1},\dots,a_{t}=\operatorname{F}(R)\}$ where $a_{i}<a_{i+1}$ for $i=1,\dots,t-1$. By Theorem 2.4 in [10], $R$ is almost symmetric if and only if $a_{i}+a_{t-i}=a_{t}$ for all $i\in[1,t-1]\cap\mathbb{Z}$.
@@
If $a\in[0,a_{1})$, then $a\in\mathbb{Z}\setminus R^{\prime}$ implies $a\in\mathbb{Z}\setminus R$ and $a\notin\operatorname{PF}(R)$. Since $R$ is almost symmetric, $\operatorname{F}(R)-a\in R$, and thus $\operatorname{F}(R)-a=r_{j}$ for some $j\in[k+1,n-1]\cap\mathbb{Z}\}$. This means $\operatorname{F}(R)-a-t+1=r_{j}-t+1\in R^{\prime}$, which implies $\operatorname{F}(R^{\prime})-a=r_{j}^{\prime}\in R^{\prime}$.
@@
If $a\in[a_{1},\operatorname{F}(R^{\prime})]$, then $a+t-1\in[a_{1}+t-1,\operatorname{F}(R^{\prime})+t-1]$, and so $a+t-1\in(a_{t-1},\operatorname{F}(R)]$. Thus, $a+t-1\in\mathbb{Z}\setminus R$ and $a+t-1\notin\operatorname{PF}(R)$. Since $R$ is almost symmetric, $\operatorname{F}(R)-a-t+1\in R$, and $\operatorname{F}(R)-a-t+1=r_{j}$ for some $j\in[1,k]\cap\mathbb{Z}$. This means $\operatorname{F}(R)-t+1-a=\operatorname{F}(R^{\prime})-a=r_{j}$.

### v2

@@ 1. Introduction
Throughout, we denote by $\mathbb{Z}$ and $\mathbb{N}$ the sets of integers and positive integers, respectively, and define $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. A numerical set $R$ is a subset of $\mathbb{N}_{0}$ that contains 0 and has a finite complement in $\mathbb{N}_{0}$. In particular, $\mathbb{N}_{0}$ itself is a numerical set with an empty complement. If a numerical set $R$ is closed under addition, i.e. if $x,y\in R$ implies $x+y\in R$, then it is referred to as a numerical semigroup.
@@ 2. Numerical Sets and Semigroups
If a numerical set $R$ is not equal to $\mathbb{N}_{0}$, it is referred to as proper. For a proper numerical set $R$, the complement $\operatorname{G}(R)=\mathbb{N}_{0}\setminus R$ is called the set of gaps of $R$. The genus of $R$, denoted by $\operatorname{g}(R)$, is the number of gaps of $R$, i.e. the cardinality of $\operatorname{G}(R)$. The largest element of $\operatorname{G}(R)$ is known as the Frobenius number of $R$, denoted by $\operatorname{F}(R)$. The quantity $\operatorname{C}(R)=\operatorname{F}(R)+1$ is referred to as the conductor of $R$, which satisfies that $z\in\mathbb{N}_{0}$ and $z>\operatorname{C}(R)$ implies $z\in R$. Also, we define the Frobenius of $\mathbb{N}_{0}$ to be 1-1, and its conductor to be 0.
@@
A numerical set $R$ is said to be symmetric if for all $a\in[0,\operatorname{F}(R)]\cap\mathbb{N}_{0}$, exactly one of the elements $a$ or $\operatorname{F}(R)-a$ belongs to $R$.
@@
A numerical semigroup $R$ is symmetric if it is symmetric as a numerical set, or equivalently if it has the property that $2\operatorname{g}(R)=\operatorname{F}(R)+1$. And $R$ is called pseudo-symmetric if and only if $\operatorname{F}(R)$ is even and for every $z\in\mathbb{Z}\setminus R$, either $z\in\{\frac{\operatorname{F}(R)}{2},\operatorname{F}(R)\}$ or $\operatorname{F}(R)-z\in R$. Equivalently $R$ is pseudo-symmetric if and only if $2\operatorname{g}(R)=\operatorname{F}(R)+2$.
@@ 3. Numerical Sets and Young Diagrams
Let $Y$ be a Young diagram with $n$ columns and $g$ rows. The number of boxes in a given column (respectively, row) is referred to as the length of that column (respectively, row). The hook of a box in $Y$ consists of the box itself, all boxes to its right in the same row, and all boxes below it in the same column. The total number of boxes in the hook is called the hook length of that box. For more intuition on Young diagrams, we refer to [5].
@@
[…]
Let $R=\{0,r_{1},\dots,r_{n-1},r_{n},\rightarrow\}$ be a numerical set with gap set $\operatorname{G}(R)=\{a_{1},\dots,a_{g}\}$ and associated Young diagram $Y_{R}$. By construction, $Y_{R}$ has $g$ rows and $n$ columns. From left to right, each $i$th small element of $R$ corresponds to the $(i+1)$th column of $Y_{R}$, so that $r_{0}$ corresponds to the first column, $r_{1}$ to the second column, and so on. Similarly, from bottom to top, for each gap $a_{j}$ of $R$ corresponds to the $j$th row of $Y_{R}$.
The hook length of the box in the $i$th row (from bottom to top) and the first column of $Y_{R}$ is the $i$th gap $a_{i}$ of $R$. Moreover, for each $i\in[0,n-1]\cap\mathbb{N}_{0}$, the $i$th column of $Y_{R}$ can be identified with the set of hook lengths of the boxes in it. We denote this set by $\operatorname{G}_{i}(R)$.
@@
[14, Section 2] Let $R=\{0,r_{1},\dots,r_{n-1},r_{n},\rightarrow\}$ be a numerical set with the associated Young diagram $Y_{R}$. Then:
@@
    For each $i\in[0,n-1]\cap\mathbb{N}_{0}$, we have G_i(R)={ a-r_i a∈G(R) and a¿r_i }=G(R-r_i) where $R-r_{i}=\{0,r_{i+1}-r_{i},\dots,r_{n}-r_{i},\rightarrow\}$.
@@
    $R$ is a numerical semigroup if and only if $\operatorname{G}_{i}(R)\subseteq\operatorname{G}_{0}(R)$ for all $i\in[0,n-1]\cap\mathbb{N}_{0}$, or equivalently $\bigcup_{i=0}^{n-1}\operatorname{G}_{i}(R)=\operatorname{G}(R)$.
By Proposition 3.2, it follows that $\operatorname{G}_{0}(R)=\operatorname{G}(R)$, and for each $i\in[0,n-1]\cap\mathbb{N}_{0}$, the hook length of the box in the $i$th column (from left to right) and the top row of $Y_{R}$ is $\operatorname{F}(R)-r_{i}$.
@@
For a positive integer $N$, a partition $\nu$ of $N$ is a finite non-increasing sequence of positive integers $\nu_{1}\geq\nu_{2}\geq\dots\geq\nu_{n}$ such that $\nu_{1}+\nu_{2}+\dots+\nu_{n}=N$, denoted by $\nu=[\nu_{1},\nu_{2},\dots,\nu_{n}]$. Each $\nu_{i}$ is called a part of the partition and the number $n$ of parts is referred to as the length of the partition.
In the context of Young diagrams, listing all the lengths of all rows from top to down yields a partition. Conversely, every partition $\nu=[\nu_{1},\nu_{2},\dots,\nu_{n}]$ corresponds to a Young diagram with $\nu_{1}$ columns and $n$ rows, where the length of $i$th row (from top to down) is $\nu_{i}$. This correspondence defines a bijective mapping between the set of partitions and the set of Young diagrams. Hence, we can identify a Young diagram $Y$ as $Y=\nu$ if its corresponding partition is $\nu$. For instance, the associated Young diagram $Y_{R}$ of the numerical set $R$ given in 2.1 is corresponding to the partition $[6,4,2,1,1,1]$ of $15$, and so we write $Y_{R}=[6,4,2,1,1,1]$.
Next, we present algorithms for visualizing numerical sets using Young diagrams. To facilitate their description, we begin by establishing some notation. Let $L=[l_{1},\dots,l_{n}]$ be a list of length $n$. We denote by $L[i]$ the $i$th element of $L$. For an integer $a$, we define $a\pm L$ as the list $a\pm l_{1},\dots,a\pm l_{n}]$. Additionally, $\operatorname{Filtered}(L,m\rightarrow m>0)$ denotes the list containing only the positive elements of $L$, preserving their original order.
@@
0: A list of integers starting from zero in increasing order $r_{0}=0,r_{1},\dots,r_{n}]$.
@@
0: A list of non-increasing positive integers $p_{1},\dots,p_{n}]$.
@@
0: A list of integers starting from zero in increasing order $r_{0}=0,r_{1},\dots,r_{n}]$.
@@ 4. Operations on Young Diagrams
$\underbrace{\hskip 35.3338pt}_{\text{first case}}$ $\underbrace{\hskip 35.3338pt}_{\text{second case}}$ $\underbrace{\hskip 35.3338pt}_{\text{third case}}$
@@
    The bonded sum of $Y$ and $X$, $Y\boxplus_{B}X$, is defined by glueing $X$ above $Y$ in such a way that the bottom row of $X$ is placed to right of the top row of $Y$. The resulting diagram is a Young diagram with $n+m$ columns and g+k1g+k-1 rows.
@@
    The conjoint sum of $Y$ and $X$, $Y\boxplus_{C}X$, is defined by glueing $X$ above $Y$ such that the most-left column of $X$ is placed directly on top of the right-most column of $Y$. The resulting diagram is a Young diagram with n+m1n+m-1 columns and $g+k$ rows.
@@
    $\boxplus_{B}$ == $\leftarrow$ $\rightarrow$
    $\underbrace{\hskip 47.11166pt}_{\displaystyle Y}$ $\underbrace{\hskip 47.11166pt}_{\displaystyle X}$ $\underbrace{\hskip 94.22345pt}_{\displaystyle Y\boxplus_{B}X}$
@@
    $\underbrace{\hskip 47.11166pt}_{\displaystyle Y}$ $\underbrace{\hskip 47.11166pt}_{\displaystyle X}$ $\underbrace{\hskip 94.22345pt}_{\displaystyle Y\boxplus_{E}X}$
@@
    $\boxplus_{C}$ == $\downarrow$ $\uparrow$
    $\underbrace{\hskip 47.11166pt}_{\displaystyle Y}$ $\underbrace{\hskip 47.11166pt}_{\displaystyle X}$ $\underbrace{\hskip 82.44548pt}_{\displaystyle Y\boxplus_{C}X}$
@@
Over and above that we get the conductor of the resulting numerical set $R\boxplus_{C}S$ as $\operatorname{C}(R\boxplus_{C}S)=\operatorname{C}(R)+\operatorname{C}(S)-1$, and the Frobenius of $R\boxplus_{C}S$ as $\operatorname{F}(R\boxplus_{C}S)=\operatorname{F}(R)+\operatorname{F}(S)$. By analogous descriptions conjoint and end-to-end sums of $R$ and $S$ can be formulated as well.
@@
Let $R=\{0,r_{1},\dots,r_{n}=\operatorname{C}(R),\rightarrow\}$ and $S=\{0,s_{1},\dots,s_{m}=\operatorname{C}(S),\rightarrow\}$ be numerical sets. Then we define
@@
Let $R=\{0,r_{1},\dots,r_{n},\rightarrow\}$ and $S=\{0,s_{1},\dots,s_{m},\rightarrow\}$ be numerical sets with $\operatorname{G}(R)=\{a_{1},\dots,a_{g}\}$ and $\operatorname{G}(S)=\{b_{1},\dots,b_{k}\}$, respectively. Then
@@
    $\operatorname{G}(R\boxplus_{B}S)=\{a_{1},\dots,a_{g-1},a_{g}+b_{1},\dots,a_{g}+b_{k}\}$ and $\operatorname{F}(R\boxplus_{B}S)=a_{g}+b_{k}$,
@@
Let $R=\{0,r_{1},\dots,r_{n},\rightarrow\}$ be a numerical set. Then $\mathbb{N}_{0}\boxplus_{B}R=\{-1,0,r_{1}-1,\dots,r_{n}-1,\rightarrow\}$ is not a numerical set, and $\mathbb{N}_{0}\boxplus_{C}R=\{r_{1}-1,\dots,r_{n}-1,\rightarrow\}$ is not also a numerical set unless $r_{1}=1$. However, we naturally have the following.
@@
Let $R$ and $S$ be numerical sets. Then
@@
Let $R=\{0,r_{1},\dots,r_{n},\rightarrow\}$ with $\operatorname{G}(R)=\{a_{1},\dots,a_{g}\}$ and $S=\{0,s_{1},\dots,s_{m},\rightarrow\}$ with $\operatorname{G}(S)=\{b_{1},\dots,b_{k}\}$. Then
@@ 5. Decompositions of Almost Symmetric Numerical Semigroups
[…]
@@
In case $\frac{\operatorname{C}(R)}{2}\notin R$, we select $S=\{0,r_{1},\dots,r_{k-1},r_{k}+1,\rightarrow\}$ with the gap set $\operatorname{G}(S)=\{a_{1},\dots,a_{n-k},r_{k}\}$. Then we get $\operatorname{F}(R)-\frac{\operatorname{C}(R)}{2}\in R\implies\frac{\operatorname{C}(R)}{2}-1\in R\implies\operatorname{F}(S)=r_{k}=\frac{\operatorname{C}(R)}{2}-1$ and $\operatorname{C}(S)=r_{k}+1=\frac{\operatorname{C}(R)}{2}$. The dual of $S$ is then $S^{*}=\{0,r_{k}-a_{n-k-1},\dots,r_{k}-a_{1},r_{k}+1,\rightarrow\}.$ Therefore,
@@
$\boxplus_{C}$ $\boxplus_{B}$ ==
$\underbrace{\hskip 23.55583pt}_{\displaystyle Y_{S}}$ $\underbrace{\hskip 11.77797pt}_{\displaystyle Y_{\{0,2,\rightarrow\}}}$ $\underbrace{\hskip 35.3338pt}_{\displaystyle Y_{S^{*}}}$ $\underbrace{\hskip 58.88965pt}_{\displaystyle Y_{R}=Y_{S}\boxplus_{C}Y_{\{0,2,\rightarrow\}}\boxplus_{B}Y_{S^{*}}}$
@@
$\boxplus_{B}$ $\boxplus_{C}$ ==
$\underbrace{\hskip 23.55583pt}_{\displaystyle Y_{S}}$ $\underbrace{\hskip 11.77797pt}_{\displaystyle Y_{\{0,2,\rightarrow\}}}$ $\underbrace{\hskip 47.11166pt}_{\displaystyle Y_{S^{*}}}$ $\underbrace{\hskip 70.66762pt}_{\displaystyle Y_{R}=Y_{S}\boxplus_{B}Y_{\{0,2,\rightarrow\}}\boxplus_{C}Y_{S^{*}}}$
When decomposing a symmetric numerical semigroup $R=\{0,r_{1}\dots,r_{n},\rightarrow\}$ into a numerical semigroup and its dual, some certain considerations must be made. By Theorem 5.1, when $\frac{\operatorname{C}(R)}{2}\notin R$, the natural choice would be $S=\{0,r_{1},\dots,r_{k-1},r_{k}+1,\rightarrow\}$ where $r_{k}=\frac{\operatorname{C}(R)}{2}-1$ for some $k\in[1,n]\cap\mathbb{N}_{0}$. In this case, if $r_{k}$ is not a minimal generator of $R$, then $S$ fails to be a numerical semigroup. Thus, to ensure $S$ remains a numerical semigroup, we need to consider a slightly different $S$ and a different sum.
Hence, when $R=\{0,r_{1}\dots,r_{n},\rightarrow\}$ is a symmetric numerical semigroup with $\frac{\operatorname{C}(R)}{2}\notin R$, to ensure $S$ remains a numerical semigroup in the decomposition, we need to choose that $S=\{0,r_{1},\dots,r_{k},r_{k}+2\rightarrow\}$ with $r_{k}=\frac{\operatorname{C}(R)}{2}-1$ for some $k\in[1,n]\cap\mathbb{N}_{0}$. In this way, we guarantee that $S$ is a numerical semigroup. However, the decomposition takes the form $R=S\boxplus_{C}(S^{*}-1)$ where $S^{*}-1=\{s-1\mid s\in S^{*}\cap\mathbb{N}\}\cup\{0\}$. To prove this statement, we shall use a similar way as in the proof of Theorem 5.1.
@@
[13, cf. Theorem 4.4] Let $R$ be a symmetric numerical semigroup. Then there exists a unique numerical semigroup $S$ such that
@@
Let $R$ be a numerical semigroup with type $\operatorname{t}(R)=t$, and let $\operatorname{PF}(R)=\{a_{1},\dots,a_{t}=\operatorname{F}(R)\}$ where $a_{i}<a_{i+1}$ for $i=1,\dots,t-1$. By Theorem 2.4 in [10], $R$ is almost symmetric if and only if $a_{i}+a_{t-i}=a_{t}$ for all $i\in[1,t-1]\cap\mathbb{N}_{0}$.
@@
If $a\in[0,a_{1})$, then $a\in\mathbb{Z}\setminus R^{\prime}$ implies $a\in\mathbb{Z}\setminus R$ and $a\notin\operatorname{PF}(R)$. Since $R$ is almost symmetric, $\operatorname{F}(R)-a\in R$, and thus $\operatorname{F}(R)-a=r_{j}$ for some $j\in[k+1,n-1]\cap\mathbb{N}$. This means $\operatorname{F}(R)-a-t+1=r_{j}-t+1\in R^{\prime}$, which implies $\operatorname{F}(R^{\prime})-a=r_{j}^{\prime}\in R^{\prime}$.
@@
If $a\in[a_{1},\operatorname{F}(R^{\prime})]$, then $a+t-1\in[a_{1}+t-1,\operatorname{F}(R^{\prime})+t-1]$, and so $a+t-1\in(a_{t-1},\operatorname{F}(R)]$. Thus, $a+t-1\in\mathbb{Z}\setminus R$ and $a+t-1\notin\operatorname{PF}(R)$. Since $R$ is almost symmetric, $\operatorname{F}(R)-a-t+1\in R$, and $\operatorname{F}(R)-a-t+1=r_{j}$ for some $j\in[1,k]\cap\mathbb{N}$. This means $\operatorname{F}(R)-t+1-a=\operatorname{F}(R^{\prime})-a=r_{j}$.